OpenStudy (anonymous):

Michelle rents a movie for a flat fee of $1.50 plus an additional $1.25 for each night she keeps the movie. Choose the cost function that represents this scenario if x equals the number of nights Michelle has the movie. c(x) = 1.50 + 1.25x c(x) = 1.50x + 1.25 c(x) = 2.75 c(x) = (1.50 + 1.25)x

OpenStudy (anonymous):

Okay. Well no matter what she'll be paying $1.50 since that's the flat fee. So it's $1.50 plus how much she's paying for whatever time she keeps it. If I'm paying 2 dollars a night and I'm keeping the movie for 3 nights, I would multiply 2 and 3 to find out I'm paying $6. So if x is the amount of days she is keeping the movie, you're going to multiply that by the cost per night-$1.25. So the answer is c(x)= 1.50 + 1.25x, your first choice. Do you understand?

OpenStudy (anonymous):

yes! thanks so much!!!!!

OpenStudy (anonymous):

You're welcome. :)

OpenStudy (anonymous):

hey! @JakeV8

OpenStudy (anonymous):

Let's try this one... Laura rents a movie for a flat fee of $2.00 plus an additional $0.50 for each night she keeps the movie. If Laura returned the movie and paid $5.00, how many nights did she keep the movie?

OpenStudy (anonymous):

Do you understand the idea of the cost function? If not, we can back up one step.

OpenStudy (anonymous):

# nights flat fee 0.50 per-night cost total cost ------- ------ ----------------- -------- 0 $2 0 $2 1 $2 0.50 * 1 = 0.50 $2.50 2 $2 0.50 * 2 = 1.00 $3.00 3 $2 0.50 * 3 = 1.50 $3.50 in general: x nights $2 0.50 * x c(x) = 2 + 0.50x